OpenStudy (anonymous):

Order 1 for 12 bushes and 15 trees that will cost the customer $600 Order 2 for 8 bushes and 20 trees that will cost the customer $680 What is the cost of one bush?

OpenStudy (texaschic101):

12b + 15t = 600 reduce by dividing by 3 giving us : 4b + 5t = 200 8b + 20t = 680 reduce by dividing by 4 giving us : 2b + 5t = 170 It is not necessary to reduce the equations to fidn the answer, it is just easier when you working with smaller numbers. So now we have : 4b + 5t = 200 2b + 5t = 170 --->(-1) ----------- 4b + 5t = 200 -2b - 5t = -170 (result of multiplying by -1) ------------add 2b = 30 b = 15 I know you do not need to find the trees, but I am going to just so I can check my answer. 4b + 5t = 200 4(15) + 5t = 200 60 + 5t = 200 5t = 140 t = 28 check.. 2b + 5t = 170 2(15) + 5(28) = 170 30 + 140 = 170 170 = 170 (correct) so your bushes (b) cost $ 15 a piece and your trees (t) cost $ 28 a piece